# Manifesto from Dagstuhl Perspectives Workshop 24352—Conversational Agents: A Framework for Evaluation (CAFE)

### v1

@@ Abstract
This workshop brought together $2222$ experts from both academia and industry in neighboring areas—namely Information Retrieval (IR), Recommender Systems (RS), and Natural Language Processing (NLP)—in order to envision a new framework for the evaluation of CONversational Information ACcess (CONIAC) systems and to discuss the related research challenges.
@@ 1 Introduction
The very idea of some conversational interaction with information access systems traces back several decades. For example, Codd [28] early imagined conversations as a key way to allow casual users to interact with databases in a natural way, avoiding the complexity of a formal query language. Conversations figured out by Codd [28] were very close to the idea of conversation we are still talking about nowadays, e.g., Codd [28, p. 182]:
@@
Recent years have witnessed a spurring of research on conversational agents, in multiple intersecting domains, such as Artificial Intelligence (AI) [45], Information Seeking (IS) [128], Information Retrieval (IR) [66, 7, 67], Recommender Systems (RS) [60], Natural Language Processing (NLP) [95, 78], and even more with the rise of Large Language Models (LLMs) and Retrieval-Augmented Generation (RAG) [5].
@@
In order to answer this question, in Fig. 1, we present an example of how we imagine future conversations between humans and agents, inspired by Di Noia et al. [37]. We choose product search, a well-explored domain where it should be easier to answer “yes, it is a solved problem!” Differently from the early conversation imagined by Codd [28] and confined to the realm of relational databases, in our case, the conversation naturally spans different types of systems and methodologies—e.g., IR systems and RSs—even if some commonalities are there, such as the system asking for clarifications when needed.
In the example, the user wants to buy a new swimsuit, but he is initially unsure or unaware of his exact requirements for this swimsuit. To support his decision-making process, the user initiates a dialog with the system: “I would like to buy a new swimming suit.” The system understands that the user’s information need is best served by a RS component. Therefore, it responds with an attribute-related question: “Great! Do you prefer elastane or polyester?” Since the user does not know the pros and cons of the two materials, he asks for an explanation of the system: “I don’t really know. What are the advantages of each material?” Initially, the system relies on a combination of IR and generation, e.g., a RAG component, to answer the user information need; afterward, it relies on a personalized RS to suggest going for polyester, since this material best matches with the user preference for more durable items. The user appreciates the suggestion and goes on asking “Which do you think are better, briefs or shorts?” Since the system does not correctly understand the term “briefs”, it relies on a clarification component to ask for an explanation, “What do you mean by briefs?” Once the user clarifies that he actually means “speedo”, the system relies on a personalized RS component to suggest some possible shorts—a type of swimsuit that better fits with the user preferences and profile—and, finally, the user picks one of them, ending the conversation. Fig. 1 also shows that, while we typically think of conversations made up just of textual utterances, we should generally consider multimodal and multimedia conversations; in this simple example, this fact is represented by images of swimsuits in some utterances.
@@ 2 CONversational Information ACcess (CONIAC)
Figure 4: The CONIAC World Model.
Our broad view encompasses aspects of conversational information access systems as defined by Balog [11]:
@@
and conversational information seeking as defined by Zamani et al. [128]:
@@ 2.1 The CONIAC Process Layer
We use the term CONIAC Process Layer, or conversational process for short, to refer to any interaction between a CONIAC agent (the system) and a user (the human). Some examples of conversational processes and mechanisms that influence the actual moves, utterances, or turns being generated are those that establish and maintain interpersonal rapport; those that organize the interaction channel through turn-taking cues, meta-communicative turns and even non-verbal actions such as gestures, facial expressions, and laughter; those that structure the informational flow of the conversation and narration as it proceeds, manage continuity and coherence, and negotiate topical shifts or switches when necessary; and those that define, modify, and address discourse referents and their properties. These include turns such as plain utterances, clarification questions, greetings, repairs, and verifications [52, 48, 53, 117]. Many conversational mechanisms are intuitively acquired by humans without instruction, though some are more explicitly trained in education or professional contexts.
More specifically, and more related to information access, conversational processes of interest include system and user collaborating in the expression of an information need, revealing the system’s capabilities and collection qualities, clarification questions using mixed initiative, and persistence of discourse referents, allowing both parties to refer back to recently mentioned items or sets; and the ability to reason about the qualities and utility of a set and to further refine it for the purpose at hand [94].
@@
Specifically, a conversational process comprises different conversational events that occur over continuous time. We mean event in the most abstract sense, ranging from simple and atomic textual utterances to complex and compound multimedia and multimodal interactions. We distinguish four conversational events—start, end, user, and system events—, that will be described in greater detail in the following paragraphs.
@@ 2.2 The CONIAC System Layer
Unlike general interactive information systems, a CONIAC system must deal with the conversational dynamics described above, in particular by taking or relinquishing the initiative when necessary to keep the conversation flowing naturally [117]. In this regard, CONIAC systems differ from many other systems because the latter often have one-shot interactions or repeated interactions that can be viewed in isolation. An example of the latter could be repeated query reformulation when using a search engine, e.g., “Hillary Clinton”, followed by “Hillary Clinton state secretary”, where each time the IR system responds with a list of documents relevant to the current query and there is no notion of “state” from one query to the next. In contrast, in our process-oriented view of CONIAC systems, this notion of state is essential, as it allows the system to use knowledge of past interactions to refine current and future interactions.
The bottom part of Fig. 4 depicts the CONIAC System Layer. We make abstraction of real-world CONIAC systems and only assume that the system will emit events that are generated by one or more components of the system, e.g., IR or RS components, essentially blurring the boundaries among components. These events can be, for example, responses to information needs, suggested items, explanations, or requests for clarification from the user. As mentioned previously, another important assumption we make as a result of our view of a conversation as a process requires that there is some shared state across interactions between the user and the system, but also between different components—IR, RS, Question Answering (QA)—that together make up a CONIAC system. A state may even be recorded across different sessions—conversational processes—with the system, resulting in personalized interactions tailored to the user’s (interaction) preferences.
@@
As an example of a possible concrete instantiation of a CONIAC system, Boutilier [18] defines a preference elicitation recommendation conversation as a Partially Observable Markov Decision Process (POMDP) model, which additionally requires a reward model to encapsulate the agent’s interaction objectives. One advantage of such a decision-theoretic system model (which may be optimized through reinforcement learning methods when exact model details are unknown) is that it provides algorithmic and learning methodologies for optimizing system actions concerning the system design. This stands in contrast to many existing conversational dialog systems leveraging rule-based decision-making that may not optimize an explicitly defined system objective.
@@ 2.3 Limitations of Current Systems
Traditional IR systems are particularly limited in their capacity to handle these scenarios because although the queries are ambiguous, the assumption in most cases is that users’ information needs are relatively well-defined from the outset. These systems are primarily designed to interpret and respond to explicit queries, focusing on accurately understanding and estimating a pre-existing, well-formed information need [25, 22, 8]. Consequently, they are less effective in situations when a user’s needs are still evolving or are only partially articulated, which is common in many conversational contexts. In such scenarios, the rigid query-response paradigm of conventional IR systems falls short, as it does not adequately support the dynamic, iterative process of refining and clarifying user intent that is essential for effective conversational information access.
Interactive Information Retrieval (IIR) systems are designed to support and even encourage user exploration within the information-seeking process, yet they are fundamentally limited by the fact that interaction is always initiated by the user. In these systems, the system consistently assumes the role of a passive responder, reacting to user queries rather than actively contributing to the dialogue. While such systems can effectively address well-defined user intents, they often fall short in scenarios where the user may have a clear goal but lacks awareness of all relevant aspects, possibilities, and available design or choice options [123, 83].
This limitation means that users might not fully explore the breadth of information or decision-making options available to them, as the system does not proactively guide them toward uncovering unknown preferences or considerations. As a result, the interaction remains user-driven, potentially leading to missed opportunities for discovery and a less comprehensive understanding of the subject matter. To overcome these challenges, more advanced systems that can take on a more active role in the interaction, such as by prompting users with relevant questions, suggesting alternative pathways, or highlighting overlooked options, are needed. These systems would better support users in navigating complex information landscapes and making more informed decisions [65, 15, 123, 83].
@@
Recommender Systems (RS), while highly effective in providing personalized content and suggestions based on user preferences, are also limited when applied to conversational scenarios. RS are inherently push-based systems, where suggestions are proactively presented to the user based on a pre-existing model of user behavior, preferences, and past interactions. Although they are beneficial in providing insights about choices, aspects, and actions that are absent in IR systems, recommender systems often lack query understanding capabilities. They do not interactively interpret or refine user intent based on evolving dialogue. Finally, both the IR and the RS approaches are mostly one-shot in the sense that they are optimized for single interaction settings. Even in session-based search [33] reformulations to an initial query result to a one-shot retrieval at the end of a session, and do not allow for intent modification or asynchronous change in intent.
@@ 3.1 The Evaluation Layer
The vision of the CONIAC World Model depicted in the previous section calls for a paradigm shift in evaluation. As shown in Fig. 5, we can imagine that the CONIAC Evaluation Layer sits on top of the CONIAC process and system layers, to allow for their assessment.
@@
The literature is rich with examples of how to approach the evaluation problem from multiple perspectives [75, 81, 58, 80, 92]. However, in general, these assessment strategies focus on a more static approach, focused mostly on overall performance with the conversational system or user satisfaction with (their interactions with) said system.
We argue that when it comes to CONIAC, evaluation should be inherently dynamic and multi-faceted. Indeed, we are aware that there is no optimal set of measurements and measures that can, for every CONIAC use case, comprehensively monitor the system (Section 2.2), the users (Section 2.1), and the constantly evolving exchanges that take place. Consider an example of a direct answer to an unambiguous question ‘what is the weather like’. This direct answer might be assessed in terms of fast-time response and correctness (e.g., the system can provide the right information for the right location). In contrast, for a more complex inquiry (e.g., a child explores to understand better a new concept (‘Can animals sleep upside down’), assessment should be more focused not just on the ability of the system to understand the information needed and provide a relevant response, but also on other perspectives inherent to this particular user. In this example, perspectives could include the ability of the system to provide suitable information and to respond in a manner that a child can comprehend [75, 6]. However, a clear pattern emerges: regardless of ‘what’ (i.e., criteria) is assessed or ‘how’ (i.e., measures), assessment of CONIAC systems cannot be done in a summative manner, based on a single measurement.
@@ 3.2 Evaluation Framework
The following Sections from 4 to 9 will go into the details of each of the items above.
@@ 3.3 Limitations of Current Conversational Evaluation
The literature is rich in algorithms and strategies for handling multiple scenarios from conversation information access. While most of them are inspired by conversational search and conversational IR literature, we have seen an interest in conversational recommendations over the last few years. Going through these works, a key commonality emerges: the lack of consensus on evaluating a conversational system in a manner that gives a holistic view of its performance and enables comparing and contrasting with other emerging technologies. Indeed, TREC tracks like TREC CAsT [33, 32, 34, 88] and TREC iKAT [3, 4] offer uniformity to comparisons for conversational search, but we do so primarily with the relevance-driven focus for evaluation. Datasets such as ReDial [77], ConveRSE [56], or LLM-REDIAL [79] enable assessment of conversation for recommendations. However, as we previously stated (and illustrated with the earlier example), we emphasize that CONIAC is a complex endeavor and should be evaluated from multiple lenses [129].

### v2

# Dagstuhl Perspectives Workshop 24352—Conversational Agents: A Framework for Evaluation (CAFE): Manifesto

@@ Abstract
This is the accepted version of the “Manifesto from Dagstuhl Perspectives Workshop 24352—Conversational Agents: A Framework for Evaluation (CAFE)” (before final copyediting). This Dagstuhl Perspectives Workshop 24352 was held 25.–30. August, 2024. For details, see https://www.dagstuhl.de/24352. Cite as: Christine Bauer, Li Chen, Nicola Ferro, Norbert Fuhr, Avishek Anand, Timo Breuer, Guglielmo Faggioli, Ophir Frieder, Hideo Joho, Jussi Karlgren, Johannes Kiesel, Bart P. Knijnenburg, Aldo Lipani, Lien Michiels, Andrea Papenmeier, Maria Soledad Pera, Mark Sanderson, Scott Sanner, Benno Stein, Johanne R. Trippas, Karin Verspoor, and Martijn C. Willemsen. Conversational Agents: A Framework for Evaluation (CAFE) (Dagstuhl Perspectives Workshop 24352). In Dagstuhl Manifestos, Volume 11, Issue 1, pp. 19–67, Schloss Dagstuhl – Leibniz-Zentrum für Informatik (2025). https://doi.org/10.4230/DagMan.11.1.19
@@
This workshop brought together $22$ experts from both academia and industry in neighboring areas—namely Information Retrieval (IR), Recommender Systems (RS), and Natural Language Processing (NLP)—in order to envision a new framework for the evaluation of CONversational Information ACcess (CONIAC) systems and to discuss the related research challenges.
@@ 1 Introduction
Every day, individuals interact with a wide variety of agents: web search engines to find information, recommender systems to identify content or products to purchase, or chatbots for customer service inquiries. Given their prevalence, let us focus on conversational agents, those that mimic human conversation using either text or spoken language, and more particularly on those enabling information access. While the literature in this area is rich [ZamaniEtAl2023], the evaluation of these systems is still limited to simplistic approaches that mostly ignore the conversational nature of the human-system interaction. In this manifesto, we specially focus on information access systems [ShahBender2024, zz-WhiteShah2025-editor] which allow us to seek, access, and use relevant information.
The very idea of some conversational interaction with information access systems traces back several decades. For example, Codd1974 early imagined conversations as a key way to allow casual users to interact with databases in a natural way, avoiding the complexity of a formal query language. Conversations figured out by Codd1974 were very close to the idea of conversation we are still talking about nowadays, e.g., Codd1974:
@@
Recent years have witnessed a spurring of research on conversational agents, in multiple intersecting domains, such as Artificial Intelligence (AI) [GaoEtAl2019], Information Seeking (IS) [ZamaniEtAl2023], Information Retrieval (IR) [KeyvanHuang2023, zz-DagstuhlSeminar19461, KieselEtAl2021], Recommender Systems (RS) [JannachEtAl2022], Natural Language Processing (NLP) [RahmaniEtAl2023, LiangEtAl2023], and even more with the rise of Large Language Models (LLMs) and Retrieval-Augmented Generation (RAG) [AllanEtAl2024].
@@
In order to answer this question, in Fig. 1, we present an example of how we imagine future conversations between humans and agents, inspired by DiNoiaEtAl2024. We choose product search, a well-explored domain where it should be easier to answer “yes, it is a solved problem!” Differently from the early conversation imagined by Codd1974 and confined to the realm of relational databases, in our case, the conversation naturally spans different types of systems and methodologies—e.g., IR systems and RSs—even if some commonalities are there, such as the system asking for clarifications when needed.
In the example, the user wants to buy a new swimsuit, but he is initially unsure or unaware of his exact requirements for this swimsuit. To support his decision-making process, the user initiates a dialog with the system: “I would like to buy a new swimming suit.” The system understands that the user’s information need is best served by a RS component. Therefore, it responds with an attribute-related question: “Great! Do you prefer elastane or polyester?” Since the user does not know the pros and cons of the two materials, he asks for an explanation of the system: “I don’t really know. What are the advantages of each material?” Initially, the system relies on a combination of IR and generation, e.g., a RAG component, to answer the user information need; afterward, it relies on a personalized RS to suggest going for polyester, since this material best matches with the user preference for more durable items. The user appreciates the suggestion and goes on asking “Which do you think are better, briefs or shorts?” Since the system does not correctly understand the term “briefs”, it relies on a clarification component to ask for an explanation, “What do you mean by briefs?” Once the user clarifies that he actually means “speedo”, the system relies on a personalized RS component to suggest some possible shorts—a type of swimsuit that better fits with the user preferences and profile—and, finally, the user picks one of them, ending the conversation. Fig. 1 also shows that, while we typically think of conversations made up just of textual utterances, we should generally consider multimodal and multimedia conversations, where more complex phenomena may happen as, e.g., interruptions in case of a spoken interaction; in this simple example, this fact is represented by images of swimsuits in some utterances.
@@ 2 CONversational Information ACcess (CONIAC)
Our broad view encompasses aspects of conversational information access systems as defined by balog2021conversational:
@@
and conversational information seeking as defined by ZamaniEtAl2023:
@@ 2.1 The CONIAC Process Layer
We use the term CONIAC Process Layer, or conversational process for short, to refer to the overall conversation flow, i.e. any interaction between a CONIAC agent (the System Flow) and a human (the User Flow), via the interface of the system in a multimedia and multimodal way, in the most general setting. Some examples of conversational processes and mechanisms that influence the actual moves, utterances, or turns being generated are those that establish and maintain interpersonal rapport; those that organize the interaction channel through turn-taking cues, meta-communicative turns and even non-verbal actions such as gestures, facial expressions, and laughter; those that structure the informational flow of the conversation and narration as it proceeds, manage continuity and coherence, and negotiate topical shifts or switches when necessary; and those that define, modify, and address discourse referents and their properties. These include turns such as plain utterances, clarification questions, greetings, repairs, and verifications [hobbs1979coherence, grosz1986attention, hobbs2014towards, trippas2020towards]. Many conversational mechanisms are intuitively acquired by humans without instruction, though some are more explicitly trained in education or professional contexts.
More specifically, and more related to information access, conversational processes of interest include system and user collaborating in the expression of an information need, revealing the system’s capabilities and collection qualities, clarification questions using mixed initiative, and persistence of discourse referents, allowing both parties to refer back to recently mentioned items or sets; and the ability to reason about the qualities and utility of a set and to further refine it for the purpose at hand [radlinski2017theoretical].
@@
Specifically, a conversational process comprises different conversational events that occur over continuous time. We mean event in the most abstract sense, ranging from simple and atomic textual utterances to complex and compound multimedia and multimodal interactions, where phenomena such as interruptions, which for instance is central to spoken interaction, may happen. We distinguish four main conversational events—start, end, user, and system events—, that will be described in greater detail in the following paragraphs.
@@ 2.2 The CONIAC System Layer
Unlike general interactive information systems, a CONIAC system must deal with the conversational dynamics described above, in particular by taking or relinquishing the initiative when necessary to keep the conversation flowing naturally [trippas2020towards]. In this regard, CONIAC systems differ from many other systems because the latter often have one-shot interactions or repeated interactions that can be viewed in isolation. An example of the latter could be repeated query reformulation when using a search engine, e.g., “Hillary Clinton”, followed by “Hillary Clinton state secretary”, where each time the IR system responds with a list of documents relevant to the current query and there is no notion of “state” from one query to the next. In contrast, in our process-oriented view of CONIAC systems, this notion of state is essential, as it allows the system to use knowledge of past interactions to refine current and future interactions.
The bottom part of Fig. 4 depicts the CONIAC System Layer. We make abstraction of real-world CONIAC systems and only assume that the system will emit events that are generated by one or more components of the system, e.g., IR or RS components, essentially blurring the boundaries among components. These events can be, for example, responses to information needs, suggested items, explanations, or requests for clarification from the user. Note that by CONIAC system, we refer to both the underlying application and its interface.
As mentioned previously, another important assumption we make as a result of our view of a conversation as a process requires that there is some shared state across interactions between the user and the system, but also between different components—IR, RS, Question Answering (QA)—that together make up a CONIAC system. A state may even be recorded across different sessions—conversational processes—with the system, resulting in personalized interactions tailored to the user’s (interaction) preferences.
@@
As an example of a possible concrete instantiation of a CONIAC system, boutilier2002pomdp defines a preference elicitation recommendation conversation as a Partially Observable Markov Decision Process (POMDP) model, which additionally requires a reward model to encapsulate the agent’s interaction objectives. One advantage of such a decision-theoretic system model (which may be optimized through reinforcement learning methods when exact model details are unknown) is that it provides algorithmic and learning methodologies for optimizing system actions concerning the system design. This stands in contrast to many existing conversational dialog systems leveraging rule-based decision-making that may not optimize an explicitly defined system objective.
@@ 2.3 Limitations of Current Systems
Traditional IR systems are particularly limited in their capacity to handle these scenarios because although the queries are ambiguous, the assumption in most cases is that users’ information needs are relatively well-defined from the outset. These systems are primarily designed to interpret and respond to explicit queries, focusing on accurately understanding and estimating a pre-existing, well-formed information need [chang2020query, carpineto2012survey:query:expansion, anand2023queryunderstandingagelarge]. Consequently, they are less effective in situations when a user’s needs are still evolving or are only partially articulated, which is common in many conversational contexts [DBLP:conf/chiir/KieselBSAH19]. In such scenarios, the rigid query-response paradigm of conventional IR systems falls short, as it does not adequately support the dynamic, iterative process of refining and clarifying user intent that is essential for effective conversational information access.
Interactive Information Retrieval (IIR) systems are designed to support and even encourage user exploration within the information-seeking process, yet they are fundamentally limited by the fact that interaction is always initiated by the user. In these systems, the system consistently assumes the role of a passive responder, reacting to user queries rather than actively contributing to the dialogue. While such systems can effectively address well-defined user intents, they often fall short in scenarios where the user may have a clear goal but lacks awareness of all relevant aspects, possibilities, and available design or choice options [white2009exploratory, Marchionini2006exploratory].
This limitation means that users might not fully explore the breadth of information or decision-making options available to them, as the system does not proactively guide them toward uncovering unknown preferences or considerations. As a result, the interaction remains user-driven, potentially leading to missed opportunities for discovery and a less comprehensive understanding of the subject matter. To overcome these challenges, more advanced systems that can take on a more active role in the interaction, such as by prompting users with relevant questions, suggesting alternative pathways, or highlighting overlooked options, are needed. These systems would better support users in navigating complex information landscapes and making more informed decisions [DBLP:journals/ftir/Kelly09, Belkin2000_helpingpeople, white2009exploratory, Marchionini2006exploratory].
@@
Recommender Systems (RS), while highly effective in providing personalized content and suggestions based on user preferences, are also limited when applied to conversational scenarios. RS are inherently push-based systems, where suggestions are proactively presented to the user based on a pre-existing model of user behavior, preferences, and past interactions. Although they are beneficial in providing insights about choices, aspects, and actions that are absent in IR systems, recommender systems often lack query understanding capabilities. They do not interactively interpret or refine user intent based on evolving dialogue. Finally, both the IR and the RS approaches are mostly one-shot in the sense that they are optimized for single interaction settings. Even in session-based search [dalton2020treccast2019conversational] reformulations to an initial query result to a one-shot retrieval at the end of a session, and do not allow for intent modification or asynchronous change in intent.
@@ 3.1 The Evaluation Layer
The vision of the CONIAC World Model depicted in the previous section calls for a paradigm shift in evaluation. As shown in Fig. 5, we can imagine that the CONIAC Evaluation Layer sits on top of the CONIAC process and system layers, to allow for their assessment, where the CONIAC system is the construct under evaluation.
The literature is rich with examples of how to approach the evaluation problem from multiple perspectives [landoni2019sonny, liu2021meta, jannach2023evaluating, lipani2021doing, penha2020challenges]. However, in general, these assessment strategies focus on a more static approach, focused mostly on overall performance with the conversational system or user satisfaction with (their interactions with) said system.
We argue that when it comes to CONIAC, evaluation should be inherently dynamic and multi-faceted. Indeed, we are aware that there is no optimal set of measurements and measures that can, for every CONIAC use case, comprehensively monitor the system (Section 2.2), the users (Section 2.1), and the constantly evolving exchanges that take place. Consider an example of a direct answer to an unambiguous question ‘what is the weather like’. This direct answer might be assessed in terms of fast-time response and correctness (e.g., the system can provide the right information for the right location). In contrast, for a more complex inquiry (e.g., a child explores to understand better a new concept (‘Can animals sleep upside down’), assessment should be more focused not just on the ability of the system to understand the information needed and provide a relevant response, but also on other perspectives inherent to this particular user. In this example, perspectives could include the ability of the system to provide suitable information and to respond in a manner that a child can comprehend [landoni2019sonny, allen2023multi]. However, a clear pattern emerges: regardless of ‘what’ (i.e., criteria) is assessed or ‘how’ (i.e., measures), assessment of CONIAC systems cannot be done in a summative manner, based on a single measurement.
@@ 3.2 Evaluation Framework
We offer details for each of the items above in Section  4 to Section 9.
We also take Fig. 6 as the workflow to be adopted to evaluate CONIAC systems. We start from defining the “Stakeholders’ Goals” and progress clock-wise to “User Aspects”, ”Tasks’, ”Criteria”, ”Methodology”, and ”Measures”. All together, ‘Stakeholders’ Goals”, “User Aspects”, and ”Tasks’ constitute the context in which the CONIAC system is evaluated, whereas ”Criteria”, ”Methodology”, and ”Measures” define how the actual evaluation is to be carried out.
@@ 3.3 Limitations of Current Conversational Evaluation
The literature is rich in algorithms and strategies for handling multiple scenarios from conversation information access. While most of them are inspired by conversational search and conversational IR literature, we have seen an interest in conversational recommendations over the last few years. Going through these works, a key commonality emerges: the lack of consensus on evaluating a conversational system in a manner that gives a holistic view of its performance and enables comparing and contrasting with other emerging technologies. Indeed, TREC tracks like TREC CAsT [dalton2020treccast2019conversational, DaltonEtAl2020, dalton:2021, owoicho:2022] and TREC iKAT [aliannejadi:2023, DBLP:conf/sigir/AliannejadiAC0A24] offer uniformity to comparisons for conversational search, but do so primarily with the relevance-driven focus for evaluation. Datasets such as ReDial [ReDial], ConveRSE [ConveRSE], or LLM-REDIAL [liang-etal-2024-llm] enable assessment of conversation for recommendations. However, as we previously stated (and illustrated with the earlier example), we emphasize that CONIAC is a complex endeavor and should be evaluated from multiple lenses [zangerle2022_fevr].